OpenStudy (anonymous):

Given the function f(x) = 3(x − 3)2 + 2, indicate the shifts that will affect the location of the vertex, and explain what effect they will have. Use complete sentences. f(x+4) f(x) + 4 f(4x) 4•f(x)

OpenStudy (anonymous):

i'm confused on this can some one explain??

OpenStudy (anonymous):

@xvddedge can you help me?

OpenStudy (anonymous):

Sorry no clue.

OpenStudy (tkhunny):

f(x) = 3x^2 Vertex (0,0) f(x) = 3(x − 3)^2 Vertex (3,0) f(x) = 3(x − 3)^2 + 2 Vertex (3,2)

OpenStudy (anonymous):

???